# ContextFlow: Context-Aware Flow Matching for Trajectory Inference from Spatial Omics Data

### v2

@@ 5 Conclusion
We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data. The entropic variant of ContextFlow is theoretically grounded, which produces OT couplings that are constrained by prior knowledge, thereby promoting stability and consistency with the imposed contextual constraints. Across three diverse datasets, we showed that ContextFlow consistently outperforms state-of-the-art baselines, even in challenging Initial Value Problem sampling settings, underscoring the importance of our contextually informed priors. In addition, we demonstrated that our framework reduces the number of biologically implausible couplings and yields coherent, temporally consistent developmental trajectories while maintaining strong quantitative performance across Wasserstein, MMD, and Energy metrics. These results highlight the value of embedding biological context into generative flow models. Future work can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or by leveraging multi-marginal OT formulations to optimize temporal flows. Looking forward, ContextFlow offers a principled foundation for modeling perturbations and disease progression, potentially bridging generative power with biological interpretability.
Impact Statement
Our work contributes a new method for modeling dynamical systems from unpaired observations with structural constraints. As such, our algorithm can contribute to biomedical domains, for example, by modeling drug perturbation responses in spatially resolved tumor tissues or by analyzing complex biological dynamics with unpaired datasets. These applications could positively impact biomedical research by enabling more data-efficient modeling of longitudinal biological processes and expediting hypothesis generation about treatment effects or underlying mechanisms. Future work should emphasize broad validation of our method using experimental or clinical ground truth, and the interpretation of results with domain experts to build appropriate trust in predictions.
Appendix A Related Work
A.1 Flow Matching
Normalizing flows provide a parametric framework for characterizing transformations of a random variable into desired distributions (Papamakarios et al., 2021). These transformations can be realized through either finite (Rezende and Mohamed, 2015) or continuous compositions (Chen et al., 2018). The loss functions used in such formulations typically require computing Jacobians or integrating the flows at each forward pass, making them computationally expensive. Flow matching (FM) (Lipman et al., 2023; Albergo and Vanden-Eijnden, 2022; Liu et al., 2022) addresses this limitation by reducing the training of the velocity field to a regression problem, thereby making normalizing flows substantially more scalable. To ensure valid conditional paths at intermediate time points, samples are coupled either randomly or via optimal transport (Pooladian et al., 2023; Tong et al., 2024). Owing to this scalability, FM has been rapidly adopted across scientific domains, including biology and the life sciences (Li et al., 2025). In transcriptomics, for example, Klein et al. (2024) employed an FM backbone to approximate OT maps for drug response modeling and cross-modal translation tasks. Entropic OT formulations have also been applied to infer cellular trajectories (Tong et al., 2024; Rohbeck et al., 2025), generate imaging-based cell morphology changes (Zhang et al., 2025), and simulate spatial transcriptomics data from histology images (Huang et al., 2025). Despite these advances, existing work does not address how to meaningfully incorporate biological prior knowledge to constrain the velocity field, limiting the biological plausibility of inferred trajectories.
A.2 Optimal Transport
Omics studies frequently generate uncoupled measurements across conditions, modalities, or time points, which must be integrated into a unified representation to provide a more comprehensive view of the underlying biology. Optimal transport (OT) has recently gained popularity for this task, as it provides a geometry-based approach to couple probability distributions (Bunne et al., 2024; Klein et al., 2025). In spatial transcriptomics (ST), several OT formulations have been introduced depending on context. For instance, Zeira et al. (2022) and Liu et al. (2023) proposed PASTE and PASTE2 to align ST data from adjacent tissue slices, while DeST-OT (Halmos et al., 2025) integrates spatio-temporal slices by modeling cell growth and differentiation. Rahimi et al. (2024) developed DOT, a multi-objective OT framework for mapping features across scRNA-seq and spatially resolved assays, and Ceccarelli et al. (2025) introduced TOAST, a spatially regularized OT framework for slice alignment and annotation transfer. While these methods are primarily designed to align biological data across space, time, or modality, they do not address the problem of trajectory inference toward biologically plausible solutions, leveraging biological priors to constrain or bias the transport plan.
Appendix B Additional Definitions
B.1 Kantarovich’s OT Formulation
Kantorovich’s formulation (Peyré et al., 2019) is a classical definition of the optimal transport (OT) problem that seeks a joint coupling to move a probability measure to another that minimizes the Euclidean distance cost, corresponding to the following minimization problem with respect to the 2-Wasserstein distance:
$\pi^{*}_{\mathrm{ot}}:={\arg\!\min}_{\pi\in\Pi(q_{0},q_{1})}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\|\bm{x}_{0}-\bm{x}_{1}\|_{2}^{2}\>d\pi(\bm{x}_{0},\bm{x}_{1}),$ (14)
where $\Pi(q_{0},q_{1})$ denotes the set of joint probability measures such that the left and right marginals are $q_{0}$ and $q_{1}$. Equation 14 can be solved in a mini-batch fashion using standard solvers such as POT (Flamary et al., 2021); however, the computational complexity is cubic in batch size.
B.2 Evaluation Metrics
@@
where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples. To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model (Chen and Guestrin, 2016) trained for each dataset.

### v3

@@ 6 Conclusion
We introduced ContextFlow, a contextually aware flow matching framework that leverages spatial priors and biologically motivated constraints to learn more plausible trajectories from snapshot spatial transcriptomic data. Our experiments showed that the entropic variant of ContextFlow (CTF-H) outperforms existing baselines, even in challenging IVP sampling settings. In particular, our method effectively reduces the number of biologically implausible couplings and generates more temporally consistent developmental trajectories, while maintaining strong quantitative performance across $2$-Wasserstein, MMD, and Energy metrics. These results highlight the value of embedding proper biological context into flow-based generative models. Future work can adapt our methods to reconstruct tissues and learn spatial latent dynamics by formulating the flow in space (rather than time), or by leveraging multi-marginal OT formulations to optimize temporal flows. Looking forward, our ContextFlow framework offers a principled foundation for modeling perturbations and disease progression, potentially bridging generative power with biological interpretability.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: Our main claims focus on demonstrating the limitations of prior flow matching techniques and on our proposed solution to mitigate them. Section 3 discusses the limitations of prior techniques, Section 4 presents the proposed methodology, and Section 5 reports comprehensive experimental results against baselines to validate our claims. Additionally, we defer more analysis to Appendix C (effect of normalization in OT cost), Appendix G.1 (biological plausibility analysis), and Appendices I and J (full ablation studies).
  5. Guidelines:
    • The answer [N/A] means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A [No] or [N/A] answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Main limitations of our work include poorly specified or noisy priors, which may bias the learned dynamics. Furthermore, while our evaluation demonstrates strong statistical and biological coherence, validation using experimental or longitudinal ground truth data would strengthen confidence in the inferred trajectories.
  10. Guidelines:
    • The answer [N/A] means that the paper has no limitation while the answer [No] means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate “Limitations” section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: Yes, the formal statements of our theorems list all the necessary assumptions. Appendix D provides their detailed proofs.
  15. Guidelines:
    • The answer [N/A] means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We’ve included the information to fully describe the algorithm in Appendix E, as well as the data processing, splits, and the settings of the packages used in our studies in Section 5 and Appendix E.2.
  20. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • If the paper includes experiments, a [No] answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: The refactored source code will be released after the review process.
  25. Guidelines:
    • The answer [N/A] means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so [No] is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://neurips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We’ve included the information to fully describe the algorithm in Appendix E, as well as the data processing, splits, and the settings of the packages used in our studies in Section 5 and Appendix E.2.
  30. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We’ve provided both the mean and standard deviation statistics across random seeds for all the reported metrics, and extensive ablations in Appendices I and J.
  35. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The authors should answer [Yes] if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g., negative error rates).
    • If error bars are reported in tables or plots, the authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We’ve included information in terms of data processing, splits, settings of packages used in our studies, runtime, etc. Importantly, all our experiments were run on a single Nvidia DGX A100.
  40. Guidelines:
    • The answer [N/A] means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: We’ve read the guidelines and confirm that we’re not in violation of any code of ethics.
  45. Guidelines:
    • The answer [N/A] means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer [No] , they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [No]
  49. Justification: The paper presents a general principled methodology to propose a flow matching model for modeling spatial omics in general, rather than specific applications with direct societal implications.
  50. Guidelines:
    • The answer [N/A] means that there is no societal impact of the work performed.
    • If the authors answer [N/A] or [No] , they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate Deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pre-trained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: [N/A]
  55. Guidelines:
    • The answer [N/A] means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: [N/A]
  60. Guidelines:
    • The answer [N/A] means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: [N/A]
  65. Guidelines:
    • The answer [N/A] means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: [N/A]
  70. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: [N/A]
  75. Guidelines:
    • The answer [N/A] means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigor, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: [N/A]
  80. Guidelines:
    • The answer [N/A] means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy in the NeurIPS handbook for what should or should not be described.
Appendix A Formal Definitions of Evaluation Metrics
@@
where $n_{i}^{\text{true}},n_{i}^{\text{pred}}$ are the number of true and predicted cells of type $i$, and $N$ is the total number of samples. To determine the cell type of generated trajectories, we employ a multi-class classifier $M_{\phi}$, implemented as an XGBoost model [Chen and Guestrin, 2016] trained for each dataset.